# Inference-Time Text-to-Video Alignment with Diffusion Latent Beam Search

### v2

@@ 5 Inference-Time Text-to-Video Alignment
We normalize the rewards with min/max values so that the range is around $[0,1]$. Our experiments aim to assess: (1) scaling the search budget and computational costs for efficient resource allocation (Section 5.1); (2) evaluating alignment performance with feedback from humans and VLMs (Section 5.2); (3) assessing scalability to capable SoTA models (Section 5.3); (4) performing detailed ablations on LA steps $T^{\prime}$, and robustness to diverse and complex prompts (Section 5.4); (5) quantitative analysis on the diversity of generated video (Appendix L.7).
@@ 5.1 Scaling Search Budget and Computational Cost
Figure 5 (Left) measures the combinatorial reward calibrated to Gemini while increasing the search budget $KB\in\{1,2,4,8,16,32\}$. DLBS improves all the calibrated rewards the best as $KB$ increases (especially $KB=16,32$), while BoN and GS eventually slow down or saturate the performance in some cases. See Appendix L.1 for results with GPT calibrated reward and Appendix L.2 for the results with $KB=64$, where we still observe the improvement.
Figure 5 (Right) demonstrates the scaling trend of DLBS, proportional to the search budget, under various choices of $K$. The results show that there is an optimal balance between the number of latent $K$ and the number of beams $B$ under the same budget. For instance, as we increase the budget to $KB=16,32$, we peak around $K=4,8,16$, which is about 25–50% of the budget. This implies that balancing possession and exploration of diffusion latents in DLBS helps search for the best outputs robustly. See Appendix L.3 for further results.
We also analyze how alignment performance scales with different DLBS configurations under fixed computational budgets, using function evaluations (NFE) and wall-clock time as cost measures. As shown in Figure 6 (Left), DLBS consistently outperforms BoN and GS across all budgets, demonstrating superior efficiency in utilizing compute. Adding the LA estimator further amplifies this advantage, offering substantial performance gains with minimal overhead. In contrast, increasing the number of diffusion steps $T$ (DLBS-T; $KB=8$, $T\in\{100,200\}$) results in only marginal improvements despite the higher computational cost. Our results suggest a clear strategy for inference-time budget allocation: prioritize enabling the LA estimator and increasing the search budget $KB$ leads to substantial performance gains, while increasing the diffusion steps $T$ provides limited benefit relative to its computational cost.
@@ 5.2 Evaluation with AI and Human Feedback
As discussed in Section 4, we obtain a manageable reward function through the reward calibration, which reduces the cost for frequent evaluation queries in inference-time search. While DLBS efficiently improves the calibrated reward (Figure 5), a natural question is whether DLBS can improve an actual assessment by VLMs or humans by optimizing their calibrated rewards. We first use each calibrated reward for DLBS, then evaluate the quality using discrete scores (from 1 to 10) from Gemini or GPT-4o. Figure 7 (Left) demonstrates that DLBS maximizing calibrated rewards can improve the original preference feedback from VLMs, as we grow the search budget.
Next, we conduct an evaluation with VideoScore [54], a metric trained on human judgments that evaluates videos across five quality criteria. Figure 7 (Middle) shows that DLBS significantly improves the quantitative evaluation based on human evaluation. Lastly, we perform pairwise comparisons between DLBS-LA and GS ($KB=1$) by three human evaluators (Figure 7; Right). The results confirm that, whatever models or prompts we choose, the quality of content generated by DLBS-LA consistently outperforms that of a baseline. This emphasizes that our proposed method, integrating reward calibration and beam search in a latent space, effectively enhances perceptual video quality. See Appendix L.10 for the details.
Figure 7: (Left) DLBS on calibrated reward can improve the original preference feedback from VLMs. As in Figure 5, we use each calibrated reward (Figure 4) for DLBS and evaluate the quality with Gemini or GPT-4o. (Middle) DLBS on calibrated reward also improves another qualitative metric, the most, VideoScore [54], which is not involved in a reward calibration. (Right) The pairwise comparisons between DLBS-LA and GS ($KB=1$) by three human raters, when DLBS-LA searches for the best latents in a diffusion process.
@@ 5.3 Scaling Model Parameters and Capabilities
We scale up the base diffusion model from Latte to the latest SoTA models such as CogVideoX-5B [19] and Wan 2.1-14B [20], and evaluate if DLBS can improve the generation quality from those larger models. Since base models are more capable here, we adopt more challenging prompts from DEVIL-very-high (22 prompts) and Movie Gen Video Bench [55] (20 prompts). See Appendix F and I.2 for the details. Figure 6 (Right) shows that our methods achieve significant improvement in calibrated reward for both models. As shown in Figure 7 (Right), human evaluation also supports that our proposed method could generally work well with any text-to-video models, even with more capable models in the future. See Appendix L.9 for other results with CogVideoX and Wan 2.1.
Figure 8: (Left) Comparison among different LA steps $T^{\prime}$ on MSRVTT-test ($KB=8$). The performance improves as the number of LA steps increases. (Right) Perceptual quality comparison on a large and complex prompt set (1,003 prompts from Movie Gen Video Bench [55]). DLBS-LA can generalize to diverse prompts.
@@ 5.4 Ablation Study
Lookahead Steps for Reward Estimate  We scale up LA steps $T^{\prime}$ to obtain an accurate reward estimate. We use MSRVTT-test ($KB=8$) and Gemini reward for experiments. Figure 8 (Left) shows that as the number of LA steps increases, the performance improves more. Even $T^{\prime}=2,3$ significantly outperforms the posterior mean, which is often used in prior works [15, 11, 41]. This is because the sub-optimal performance of inference-time search comes from the approximation errors and LA estimator can notably reduce them. As shown in Figure 5, DLBS-LA $(KB=8,T^{\prime}=6)$ achieves comparable or even outperforming results with DLBS $(KB=32)$. It is quite beneficial to spend a computation to estimate the reward accurately. See Appendix L.4 for further discussions. In addition, an ablation study on diffusion steps is shown in Appendix L.5.
A Large and Complex Prompt Set  To confirm the robustness of our approach on a large and highly diverse prompt set, we compare GS ($KB{=}1$) with DLBS-LA ($KB{=}8,\;T^{\prime}{=}6$) with all the prompts in Movie Gen Video Bench [55], which comprises 1,003 prompts (Figure 8; Right). We observe that DLBS-LA consistently delivered substantially higher alignment rewards, demonstrating that DLBS generalizes effectively to complex prompt distributions.
@@ 6 Related Works
Classifier guidance [35, 56] has been the most popular to enhance text-content alignment. On top of that, recent works [11, 39] leverage reward or external feedback at inference time by selecting better latents [57], which probably achieve higher rewards during the reverse process. Kim et al. [58] propose twisted SMC [37] with reward gradient, which is not suitable for non-differentiable feedback and for domains such as video, where reward gradient needs a huge memory cost. Gradient-free methods [59, 17] such as SMC [41] or greedy search [15] often exhibit sub-optimal results affected by inaccurate reward estimates from noisy latents. Yeh et al. [60] uses ODE to estimate the reward, but it highly depends on Karras sampler [28] to avoid numerical instability. In contrast, we address the error propagation from inaccurate reward estimates with beam search and lookahead estimator via deterministic DDIM, which is more popular and stable. Our methods work more scalably when allocating more computation budget at inference time. See Appendix M for further related works.
@@ 7 Discussion and Limitation
Our reward calibration assumes that VLMs work as a proxy of human evaluation, and we demonstrate both qualitative and quantitative improvement of the video quality. It would be important to extend it to human feedback, not limited to VLMs. In practice, we often do implicit or explicit best-of-N sampling for video generation. In contrast, DLBS-LA exhibits much better computational efficiency. Spending more computation at inference time significantly improves perceptual quality, but it is orthogonal compared to speeding up the sampling process via distillation [61, 62], architecture changes [63], or parallel sampling [64]. We believe both high-quality and speedy sampling have practical needs and should be balanced.
@@ 8 Conclusion
This paper studies which metrics we should optimize and how to optimize them for better text-to-video generation. We point out that feedback from humans or capable VLMs reflects multiple dimensions of video quality, so optimizing an existing metric alone is insufficient; rather, we should calibrate the reward by combining. Our DLBS with LA estimator reduces the error propagation from the inaccurate reward estimate. We demonstrate DLBS is the most scalable, efficient and robust inference-time search that significantly improves video quality under the same computational costs. We hope our work encourages more uses of inference-time computation for text-to-video models.
@@ Acknowledgements
We thank Po-Hung Yeh, Shohei Taniguchi, Kuang-Huei Lee, Arnaud Doucet, Heiga Zen, and Yusuke Iwasawa for their support and helpful discussion on the initial idea of this work. HF was supported by JSPS KAKENHI Grant Number JP22J21582 (by March, 2025), and MS was supported by JSPS KAKENHI Grant Number JP23H04974.
@@ Appendix B Implementation Details
B.1 Code.
Our implementation for the experiment is available at https://anonymous.4open.science/r/T2V-Diffusion-Search-537B.
B.2 Models.
Our experiments cover three text-to-video diffusion models.
@@ B.3 Hyperparameters
  • CogVideoX: DDIM scheduler with the original settings and $w_{\text{cfg}}=6.0$. Owing to computational resource constraints, we limit the frame length to 17 per sample.
  • Wan 2.1: DPMSolver++ with guidance scale $w_{\text{cfg}}=5.0$. For the same computational reasons, the spatial resolution is limited to $832\times 480$ and the frame length to 33 per sample.
B.4 Hardware configuration
@@
B.5 AI-feedback endpoints
We use API endpoints: gemini-1.5-pro-002 and gpt-4o-2024-11-20.
@@ Appendix C Details of Metric Rewards
Appendix D Algorithms with Continuous-time Diffusion Process
@@
1:  $\textbf{z}^{1}_{t_{0}},\cdots,\textbf{z}^{B}_{t_{0}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$   $\triangleright$ Initial $B$ beams
@@
14:         $\triangleright$ Search $B$ higher-reward beams from $KB$ latents
@@
2:  $\tilde{s}(u)\in\{s,\dots,\lfloor\frac{M^{\prime}-u}{M^{\prime}}s+\frac{u}{M^{\prime}}M\rfloor,\dots,\lfloor\frac{1}{M^{\prime}}s+\frac{M^{\prime}-1}{M^{\prime}}M\rfloor,M\}$
@@
6:     $\tilde{\textbf{z}}_{t_{\tilde{s}(u)}|t_{\tilde{s}(u-1)}}=\frac{\sigma_{t_{\tilde{s}(u)}}}{\sigma_{t_{\tilde{s}(u-1)}}}\mathbf{z}_{t_{\tilde{s}(u-1)}}^{ij}-\alpha_{t_{\tilde{s}(u)}}(e^{-h}-1)z_{\theta}(\mathbf{z}^{ij}_{t_{\tilde{s}(u-1)}}))$
@@
8:  return: $(\textbf{z}_{t_{s}},\tilde{\textbf{z}}_{t_{M}|t_{\tilde{s}(0)}})$triangleright Latent and LA estimator
Appendix E Theoretical Analysis on Lookahead Estimator

### v3

@@ 5 Inference-Time Text-to-Video Alignment
Our experiments aim to assess: (1) scaling the search budget and computational costs for efficient resource allocation (Section 5.1); (2) evaluating alignment performance with feedback from humans and VLMs (Section 5.2); (3) assessing scalability to capable SoTA models (Section 5.3); (4) quantitative analysis on the diversity of generated video (Section 5.4); (5) validating that DLBS is complementary to fine-tuning methods (Section 5.5); (6) performing detailed ablations on LA steps $T^{\prime}$, and robustness to diverse and complex prompts (Section 5.6).
@@ 5.1 Scaling Search Budget and Computational Cost
Figure 5 (Left) measures the combinatorial reward calibrated to Gemini while increasing the search budget $KB\in\{1,2,4,8,16,32\}$. DLBS improves all the calibrated rewards the best as $KB$ increases (especially $KB=16,32$), while BoN and GS eventually slow down or saturate the performance in some cases. See Appendix M.1 for results with GPT calibrated reward and Appendix M.2 for the results with $KB=64$, where we still observe the improvement.
Figure 5 (Right) demonstrates the scaling trend of DLBS, proportional to the search budget, under various choices of $K$. The results show that there is an optimal balance between the number of latent $K$ and the number of beams $B$ under the same budget. For instance, as we increase the budget to $KB=16,32$, we peak around $K=4,8,16$, which is about 25–50% of the budget. This implies that balancing possession and exploration of diffusion latents in DLBS helps search for the best outputs robustly. See Appendix M.3 for further results.
We also analyze how alignment performance scales with different DLBS configurations under fixed computational budgets, using the number of function evaluations (NFE) and wall-clock time as cost measures. As shown in Figure 6 (Left), DLBS consistently outperforms BoN and GS across both budgets, demonstrating superior efficiency in utilizing compute. Adding the LA estimator further amplifies this advantage, offering substantial performance gains with minimal overhead. In contrast, increasing the number of diffusion steps $T$ (DLBS-T; $KB=8$, $T\in\{100,200\}$) results in only marginal improvements despite the higher computational cost. Our results suggest a clear strategy for inference-time budget allocation: prioritize enabling the LA estimator and increasing the search budget $KB$ leads to substantial performance gains, while increasing the diffusion steps $T$ provides limited benefit relative to its computational cost.
@@
Figure 7: (Left) DLBS on calibrated reward can improve the original preference feedback from VLMs. As in Figure 5, we use each calibrated reward (Figure 4) for DLBS and evaluate the quality with Gemini or GPT-4o. (Right) DLBS on calibrated reward also improves another qualitative metric, the most, VideoScore [55], which is not involved in a reward calibration.
@@ 5.2 Evaluation with AI and Human Feedback
As discussed in Section 4, we obtain a manageable reward function through the reward calibration, which reduces the cost for frequent evaluation queries in inference-time search. While DLBS efficiently improves the calibrated reward (Figure 5; Left), a natural question is whether DLBS can improve an actual assessment by VLMs or humans by optimizing their calibrated rewards. We first use each calibrated reward for DLBS, then evaluate the quality using discrete scores (from 1 to 10) from Gemini or GPT-4o. Figure 7 (Left) demonstrates that DLBS maximizing calibrated rewards can improve the original preference feedback from VLMs, as we grow the search budget.
Next, we evaluate with VideoScore [55], a metric trained on human judgments that evaluates videos across five quality criteria. Figure 7 (Right) shows that DLBS significantly improves the quantitative evaluation based on human evaluation. Lastly, we perform pairwise comparisons between DLBS-LA ($KB=8$, T=6T’=6, $\text{NFE}=2500$) and BoN ($KB=64$, $\text{NFE}=3200$) by three human evaluators (Figure 8; Left). The results confirm that, whatever models or prompts we choose, the quality of content generated by DLBS-LA consistently outperforms that of a baseline despite requiring fewer NFEs. This emphasizes that our proposed method, integrating reward calibration and beam search in a latent space, effectively enhances perceptual video quality. See Appendix M.11 for the details.
@@ 5.3 Scaling Model Parameters and Capabilities
We scale up the base diffusion model from Latte to the latest SoTA models, such as CogVideoX-5B [19] and Wan 2.1-14B [20], and evaluate if DLBS can improve the generation quality from those larger models. Note that we use SDE-DPMSolver++ [56] for Wan 2.1 experiments. Since base models are more capable here, we adopt more challenging prompts from DEVIL-very-high (22 prompts) and Movie Gen Video Bench [57] (20 prompts). See Appendix G and Appendix J.2 for the details. Figure 6 (Right) shows that our methods achieve significant improvements in calibrated reward for both models. As shown in Figure 8 (Left), human evaluation also supports that our proposed method could generally work well with any text-to-video models, even with more capable models in the future. See Appendix M.9 for results with a maximum frame length, and Appendix M.10 for further results with CogVideoX and Wan 2.1.
Figure 8: (Left) The pairwise comparisons of DLBS-LA ($KB=8$, T=6T’=6, $\text{NFE}=2500$) and BoN ($KB=64$, $\text{NFE}=3200$) by three human raters, when DLBS-LA searches effectively for the best latents in a diffusion process. (Right) Alignment–diversity tradeoff ($KB=32$). The mean pairwise distance (MPD) of ViCLIP embeddings is used as a measure of diversity. DLBS and DLBS-LA ($T^{\prime}=6$) achieve high performance while maintaining higher diversity.
5.4 Alignment-Diversity Tradeoff
Alignment for diffusion models can steer desirable outputs, but it is said that the diversity of generated samples or the performance of original models often degrade [10, 37]. While inference-time search does not change or degrade the model itself, we here compare the diversity of samples among BoN, GS, DLBS, and DLBS-LA. We measure the sample diversity as the mean pairwise distance of ViCLIP [52] embeddings (see Appendix D). Figure 8 (Right) reveals that DLBS and DLBS-LA achieve high performance with higher diversity than BoN or GS. This exhibits a benefit from the wider possession and exploration of diffusion latents in DLBS and DLBS-LA.
5.5 DLBS is Compatible with Finetuning
In image generation, Ma et al. [17] has shown that allocating additional computation at inference time can be more effective than relying solely on post-training approaches. We find a similar trend in video generation. To examine this, we apply a representative fine-tuning method, VideoDPO [58], to VideoCrafter2 [3]. As shown in Table 5.5, VideoDPO alone brings negligible improvement over the baseline. However, when combined with DLBS, the performance increases substantially on both DEVIL-high and MSRVTT-test. These results indicate that DLBS is complementary to post-training methods, enabling further performance gains even after fine-tuning. See Appendix M.7 and Appendix M.8 for results comparing DLBS with other alignment methods.
Table 1: Performance of DLBS with DPO finetuned VideoCrafter2 on DEVIL-high and MSRVTT-test datasets. While DPO alone yields marginal improvements, combining it with DLBS leads to notable gains, demonstrating the compatibility of inference-time search with fine-tuning approaches.
Method DEVIL-high MSRVTT-test
VideoCrafter2 0.337 0.555
+ DPO 0.335 0.556
+ DPO & DLBS 0.359 0.576
Figure 9: (Left) Comparison of different LA steps $T^{\prime}$ on MSRVTT-test ($KB=8$). The performance improves as the number of LA steps increases. (Right) Perceptual quality comparison on a large and complex prompt set (1,003 prompts from Movie Gen Video Bench [57]). DLBS-LA can generalize to diverse prompts.
@@ 5.6 Ablation Study
Lookahead Steps for Reward Estimate  We scale up LA steps $T^{\prime}$ to obtain an accurate reward estimate. We use MSRVTT-test ($KB=8$) and Gemini reward for experiments. Figure 9 (Left) shows that as the number of LA steps increases, the performance improves more. Even $T^{\prime}=2,3$ significantly outperforms the posterior mean, which is often used in prior works [15, 11, 41]. This is because the sub-optimal performance of inference-time search comes from the approximation errors, and LA estimator can notably reduce them. As shown in Figure 5 (Left), DLBS-LA $(KB=8,T^{\prime}=6)$ achieves comparable or even outperforming results with DLBS $(KB=32)$. It is quite beneficial to spend a computation to estimate the reward accurately. See Appendix M.4 for further discussions. In addition, an ablation study on diffusion steps is shown in Appendix M.5.
A Large and Complex Prompt Set  To confirm the robustness of our approach on a large and highly diverse prompt set, we compare GS (KB=1KB{=}1) with DLBS-LA ($KB{=}8,\;T^{\prime}{=}6$) with all the prompts in Movie Gen Video Bench [57], which comprises 1,003 prompts (Figure 9; Right). We observe that DLBS-LA consistently delivered substantially higher alignment rewards, demonstrating that DLBS generalizes effectively to complex prompt distributions. Additionally, we also assess reward transferability by applying weights calibrated on DEVIL-high and DEVIL-medium to MSRVTT-test prompts, consistently improving scores (see Appendix J.4).
@@ 6 Related Works
Classifier guidance [35, 59] has been the most popular to enhance text-content alignment. On top of that, recent works [11, 39] leverage reward or external feedback at inference time by selecting better latents [60], which probably achieve higher rewards during the reverse process. Kim et al. [61] propose twisted SMC [37] with reward gradient, which is not suitable for non-differentiable feedback and for domains such as video, where reward gradient needs a huge memory cost. Gradient-free methods [62, 17] such as SMC [41] or greedy search [15] often exhibit sub-optimal results affected by inaccurate reward estimates from noisy latents. Yeh et al. [63] uses ODE to estimate the reward, but it highly depends on Karras sampler [28] to avoid numerical instability. In contrast, we address the error propagation from inaccurate reward estimates with beam search and lookahead estimator via deterministic DDIM, which is more popular and stable. Our methods work more scalably when allocating more computation budget at inference time. See Appendix N for further related works.
@@ 7 Discussion and Limitation
Our reward calibration assumes that VLMs serve as a proxy for human evaluation, and we demonstrate both qualitative and quantitative improvements in video quality through evaluations by VLMs and human raters. In future work, incorporating more specialized and accurate evaluators (e.g., reward models that focus on physical laws [9]) could enable a more fine-grained analysis. In practice, we often do implicit or explicit best-of-N sampling for video generation. In contrast, DLBS-LA exhibits much better computational efficiency. Spending more computation at inference time significantly improves perceptual quality, but it is orthogonal compared to speeding up the sampling process via distillation [64, 65], architecture changes [66], or parallel sampling [67]. We believe both high-quality and speedy sampling have practical needs and should be balanced.
@@ 8 Conclusion
This paper studies which metrics we should optimize and how to optimize them for better text-to-video generation. We point out that feedback from humans or capable VLMs reflects multiple dimensions of video quality, so optimizing an existing metric alone is insufficient; rather, we should calibrate the reward by combining. Our DLBS with LA estimator reduces the error propagation from the inaccurate reward estimate. We demonstrate that DLBS is the most scalable, efficient, and robust inference-time search that significantly improves video quality under the same computational costs. We hope our work encourages more uses of inference-time computation for text-to-video models.
@@ Acknowledgements
We thank Po-Hung Yeh, Shohei Taniguchi, Kuang-Huei Lee, Arnaud Doucet, Heiga Zen, Robin Scheibler, and Yusuke Iwasawa for their support and helpful discussion on the initial idea of this work. HF was supported by JSPS KAKENHI Grant Number JP22J21582 (by March 2025), and MS was supported by JSPS KAKENHI Grant Number JP23H04974.
@@ Appendix B Implementation Details
Code   Our implementation for the experiments are available at https://github.com/shim0114/T2V-Diffusion-Search.
Models   Our experiments cover three text-to-video diffusion models.
@@
  • CogVideoX: DDIM scheduler with the original settings and $w_{\text{cfg}}=6.0$. Owing to computational resource constraints, we limit the frame length to 17 per sample. [^3] The experiments reported in Appendix M.9 were conducted with a different number of frames.
  • Wan 2.1: DPMSolver++ with guidance scale $w_{\text{cfg}}=5.0$. For the same computational reasons, the spatial resolution is limited to $832\times 480$ and the frame length to 33 per sample. [^3] The experiments reported in Appendix M.9 were conducted with a different number of frames.
Hardware configuration
@@
AI-feedback endpoints   We use API endpoints: gemini-1.5-pro-002 and gpt-4o-2024-11-20.
@@ Appendix C Details of Metric Rewards
Appendix D Details of Sample Diversity
We measure the sample diversity as the mean pairwise distance of ViCLIP [52] embeddings to quantify the diversity in videos, inspired by the approach for evaluating diversity in images [61]. Specifically, given $N$ generated video samples, we first extract ViCLIP embeddings $\mathbf{v}^{\text{video},(i)}$ for each sample $i$. The pairwise diversity score is then computed as the mean pairwise distance:
Here, $\langle\mathbf{v}^{\text{video},(i)},\mathbf{v}^{\text{video},(j)}\rangle$ denotes the cosine similarity between the ViCLIP embeddings of two generated videos $i$ and $j$. This formulation is similar to Equation 14, but in the case of pairwise distance computation, we take the pairwise mean of 1(cosine similarity)1-$\text${(cosine similarity)} to obtain a diversity measure.
Appendix E Algorithms with Continuous-time Diffusion Process
@@
[…]
@@
14:    $\triangleright$ Search $B$ higher-reward beams from KBKB latents
@@
[…]
@@
6:  $)\tilde{\textbf{z}}_{t_{\tilde{s}(u)}|t_{\tilde{s}(u-1)}}=\frac{\sigma_{t_{\tilde{s}(u)}}}{\sigma_{t_{\tilde{s}(u-1)}}}\mathbf{z}_{t_{\tilde{s}(u-1)}}^{ij}-\alpha_{t_{\tilde{s}(u)}}(e^{-h}-1)z_{\theta}(\mathbf{z}^{ij}_{t_{\tilde{s}(u-1)}}$))
@@
8:return: $(\textbf{z}_{t_{s}},\tilde{\textbf{z}}_{t_{M}|t_{\tilde{s}(0)}})$ $\triangleright$ Latent and LA estimator
Appendix F Theoretical Analysis on Lookahead Estimator